the length of a rectangular greenhouse is three times the width. if the perimeter of the greenhouse is 96 feet, what are the dimensions of the greenhouse?
Answer by TutorDelphia(193)

They tell you that L equals 3 times W:
l=3w
The formula for perimeter is just to add up all the sides (2 sides are the length and 2 are the width)
2l+2w=96
we can divide both sides by 2 and get
l+w=48
we know l=3w so we can sub that in
3w+w=48
combine like terms
4w=48
divide both sides by 4
w=12
l=3*12
l=36